OpenStudy (anonymous):

The function m(n) = 20(1.2)n – 1 approximates the number of members in a club n years after it was started, and the function f(n) = 15 models the yearly fee each member pays during year n. Which function approximates the total amount of money collected for yearly fees in year n? t(n) = 360^n – 1 t(n) = 300(1.2)^n – 1 <<< t(n) = 35(1.2)^n – 1 t(n) = 20(1.2)^n – 1+ 15

OpenStudy (solomonzelman):

\(\large\color{blue}{ m(n) = 20(1.2)^{n-1} }\) \(\large\color{blue}{ f(n) = 15 }\) \(\large\color{blue}{ f(n) \times m(n) }\)

OpenStudy (solomonzelman):

f(n) is a cost per person, and m(n) is a number of people. So the product of the two, would be the cost for all people in a certain year.

OpenStudy (anonymous):

@jim_thompson5910 @SolomonZelman So, I'm right? :)

OpenStudy (solomonzelman):

no.

OpenStudy (solomonzelman):

lets call what the m(n) equals, "D" So, this is what you are saying. n per D = n+D

OpenStudy (solomonzelman):

it is more like multiplication.

OpenStudy (solomonzelman):

yes, B is right (adited), smart choice...

OpenStudy (solomonzelman):

D wouldn't be right, but B is.